# Nonstandard Lagrangians and branched Hamiltonians: A brief review

### v1

@@ III.2 Hamiltonian aspects
Now that we are equipped with the Lagrangian that describes the Liénard system, we can move on to its Hamiltonian aspects. The conjugate momentum is found to be
@@
Notice that if Eq. (23) admits branching, then so does the Hamiltonian [Eq. (24)]. Below, we discuss a concrete example.
@@ III.3 A concrete example
We discuss an example now, which will help us understand the framework discussed above. In particular, we will also encounter the notion of momentum-dependent mass BGG ; brp ; ML , a concept that has generated some interest in the recent times, especially within the quantum-mechanical framework. We consider the case where $f(x)=x$ and $g(x)=x-x^{3}$ agp ; Eq. (14) is satisfied for $\ell=1,-2$. We consider each case separately now.
@@ III.3.1 Case with $\ell=1$
In the case where $\ell=1$, we have
$p=p(x,\dot{x})=\frac{1}{2}\left(\dot{x}+x^{2}-1\right)^{2}\quad\implies\quad\dot{x}=\dot{x}(x,p)=1-x^{2}\pm\sqrt{2p},$ (25)
which points towards branching. Notice that branching originates from the nonlinear dependence between $p$ and $\dot{x}$ in the equation $p=\frac{\partial L}{\partial\dot{x}}$. The corresponding branched Hamiltonians turn out to be
@@ III.3.2 Case with $\ell=-2$
implying that there is no branching, because $\dot{x}$ can be solved uniquely as a function of the momentum. A straightforward calculation reveals that the Hamiltonian turns out to be
@@
wherein, there is only one branch. In Figs. (3) and (4), we have plotted $\dot{x}=\dot{x}(x,p)$ and $H=H(x,p)$. An intriguing aspect of the Hamiltonian given in Eq. (28) is that it may be expressed as
@@
This resembles a standard Hamiltonian with the roles of coordinate and momentum being interchanged; its then tempting to interpret $m(p)=p^{-1}$ as a momentum-dependent mass. The quantization of such systems proceeds in the momentum space, often with the notion of momentum-dependent mass (see for example, Ref. BGG ).
@@
IV The $v^{4}$ model
@@
where $v$ is the velocity [^6] Now and in subsequent discussions, we will denote $\dot{x}=v$. and $\kappa>0$ is a coupling parameter. Corresponding to Eq. (30), the conjugate momentum is a cubic function in $v$ that is given by
@@
which, like $L(v)$, is also a multi-valued function (with cusps) in the conjugate momentum $p$, since each given $p$ corresponds to one or three values of $v$ as shown in Eq. (31).
Hence, for systems with a non-convex Lagrangian such as Eq. (30), the construction of the corresponding Hamiltonian in conjugate momentum variable is not unique. A similar incertitude is also encountered in cosmology models ctv ; nhm , in generalized schemes of Einstein gravity which involve topological invariants, and in theories of higher-curvature gravity ctj . We note in passing that if one is given a single-valued Lagrangian $L(x,v)$, and define it according to $L(x,v)=x^{2}-V(v)\,$, rather than the usual form $L(x,v)=v^{2}-V(x)\,$, and that if the $p$ or $v$ dependence is non-convex, then branched functions are always encountered, as a result of employing the Legendre transformation, despite having started with a single-valued Lagrangian or Hamiltonian function.
V A generalized class of Lagrangians and branched Hamiltonians
V.1 Velocity-independent potentials
@@
$p=p(v)=\left(\frac{1}{4}\right)^{\frac{2}{2k+1}}\frac{1}{(v-1)^{\frac{2}{2k+1}}}\,,$ (35)
Inverting the relation we observe at once that the velocity variable $v(p)$ emerges as a double-valued function of $p$:
@@
From a classical perspective, in order to avoid an imaginary $v(p)$, one needs to address a nonnegative $p$. This in turn implies that the slope $\frac{\partial L}{\partial v}$ is always positive. It is interesting to note that for the $k=1$ case we are led to the quantum-mechanical supersymmetric structure for the difference $H_{\pm}(x,p)-V(x)$, which reads $p\pm\frac{1}{2\sqrt{p}}$, in the momentum space. The associated spectral properties have been analyzed in the literature ruby ; BGG . We end our discussion on this example by noting that in the special case where $V(x)=x^{2}$, the branched Hamiltonian is $H_{\pm}(x,p)=x^{2}+U_{\pm}(p)$, wherein it appears as if the roles of the coordinate and the momentum have been interchanged with $U_{\pm}(p)$ being a momentum-dependent potential that exhibits two branches.
V.2 Velocity-dependent potentials
@@
Unfortunately, since a Hamiltonian has to be a function of the coordinate and its corresponding canonical momentum, the generality of the form of $H_{\pm}(x,p)$ as derived above is of little use unless we have an explicit inversion of Eq. (39) giving $v=v(p)$. We therefore have to go for the specific cases of $k$ and $U(v)$.
V.2.1 A special case
@@
where $\mu^{3/2}$ has been replaced by $\gamma$. As a final comment, the special case corresponding to $\lambda=0$ and $\gamma=1/4$ conforms to the Hamiltonian advanced in CZ [Eq. (37)].
VI Two more classes of branched Hamiltonians
VI.1 Example 1
As an extension of Eq. (33), the following higher-power Lagrangian was proposed in bspf :
@@ VI.1 Example 1
where we notice that the coefficient $\Lambda$ is non-negative for $0\leq m<\frac{1}{2}$. The main difference from Eq. (33) is in the choice of a general function $\sigma(x)$ in place of $\sigma(x)=-1$ as in Eq. (33). The other point is that the inverse exponent with respect to the model of Curtright and Zachos CZ has been taken for convenience of calculus. We have omitted the explicit potential function assuming that the interaction re-appears in a more natural manner via a suitable choice of an auxiliary free parameter $\delta$ and that of a nontrivial function $\sigma(x)$. As long as our Lagrangian $L(x,v)$ is of a nonstandard type, we will not feel disturbed by the absence of the explicit potential $V(x)$.
@@ VI.1.1 A special case
VI.2 Example 2
As a final example, let us turn to an illustration where $L(x,v)$ is of the reciprocal kind and is defined to be bspf
@@ VI.2 Example 2
It should be remarked that as $\lambda\rightarrow 0$, Eq. (55) is just the trial Lagrangian given in Eq. (5), for the choice $\mu(x)=ax^{2}$ and under a suitable identification of the constant parameters. We end by noting that Eq. (58) can be expressed with a momentum-dependent mass as
$H_{\pm}(x,p)=\frac{x^{2}}{2m(p)}+U_{\pm}(p),\quad\quad m(p)=\frac{3}{2sp},\quad\quad U_{\pm}(p)=\frac{3}{s}\lambda p\pm 2\sqrt{\frac{p}{s}}.$ (59)
VII Concluding remarks
In our present short review on the existence of nonstandard Lagrangians, we emphasized upon the associated branched Hamiltonians. Various different examples were discussed; in all of them, the velocity dependence of Lagrangian was not of (homogenous) degree two but contained either powers larger than two or negative powers. This resulted in a nonlinear relationship between the generalized velocity and the conjugate momentum, leading to a multi-valued behavior of the velocity when solved as a function of the momentum (and perhaps the coordinate).

### v2

@@ III.2 Hamiltonian aspects
With the Lagrangian that describes the Liénard system, we can move on to its Hamiltonian aspects. The conjugate momentum is found to be
@@
Notice that if (21) admits branching, then so does the Hamiltonian (22). Below, we discuss a concrete example.
@@ III.3 A concrete example
An illustrative example will help us understand the framework discussed above. In particular, this will also enable us to turn our attention to the notion of momentum-dependent mass BGG ; brp ; ML , a concept that has generated some interest in the recent times, especially within the quantum-mechanical physics-oriented model-building approaches.
We will consider the case where $f(x)=x$ and $g(x)=x-x^{3}$ agp ; (15) is satisfied for $\ell=1,-2$. For the sake of clarity, we will consider each case separately.
@@ III.3.1 Case with $\ell=1$
In the case of $\,\ell=1$ we have
$p=p(x,\dot{x})=\frac{1}{2}\left(\dot{x}+x^{2}-1\right)^{2}\quad\implies\quad\dot{x}=\dot{x}(x,p)=1-x^{2}\pm\sqrt{2p}.$ (23)
This points towards branching. Notice that branching originates from the nonlinear dependence between $p$ and $\dot{x}$ in the equation $p=\frac{\partial L}{\partial\dot{x}}$. The corresponding branched Hamiltonians turn out to be
@@ III.3.2 Case with $\ell=-2$
implying that there is no branching, because $\dot{x}$ can be extracted, uniquely, as a function of the momentum. A straightforward calculation reveals that the Hamiltonian turns out to be
@@
wherein, there is only one branch.
In Figs. (3) and (4), we have plotted $\dot{x}=\dot{x}(x,p)$ and $H=H(x,p)$. An intriguing aspect of the Hamiltonian (26) is that it may be expressed as
@@
This resembles a standard Hamiltonian, only with the roles of coordinate and momentum being interchanged.
It is then certainly tempting to interpret $m(p)=p^{-1}$ as a momentum-dependent mass. Also the quantization of such systems proceeds, in momentum space, often in the context referring to the notion of momentum-dependent mass (see for example, Ref. BGG ). Still, in our particular model such a mass becomes singular (infinite) in the zero-momentum limit. In the same limit, moreover, also the potential itself is vanishing, i.e., a consistent physical interpretation of the system would require a suitable regularization of the limiting process.
Marginally, let us add that one of the regularization recipes which appeared applicable to the quantum version of our very specific model (27) has been proposed and tested in an older paper uchac . Based on an ad hoc  complexification of the momentum and on a certain rather sophisticated strong-coupling perturbation-expansion technique the recipe has been even found to provide the numerically fairly-reliable spectra, in certain ranges of couplings at least.
@@
IV A generalized class of Lagrangians yielding branched Hamiltonians
Let us note that if one is given a single-valued Lagrangian $L(x,v)$ and defines it according to formula $L(x,v)=x^{2}-V(v)\,$ rather than according to the usual recipe $L(x,v)=v^{2}-V(x)\,$ and, moreover, if the $p$ or $v$ dependence is non-convex, then, as a result of employing the Legendre transformation, the branched functions are always encountered despite our having started from a single-valued Lagrangian or Hamiltonian function.
IV.1 The $v^{4}$ model
@@
where $v$ is the velocity [^10] Now and in subsequent discussions, we will denote $\dot{x}=v$. and $\kappa>0$ is a coupling parameter. Corresponding to (28), the conjugate momentum is a cubic function in $v$ that is given by
@@
which, like $L(v)$, is also a multi-valued function (with cusps) in the conjugate momentum $p$, since each given $p$ corresponds to one or three values of $v$ as shown in (29).
For systems with a non-convex Lagrangian as sampled by (28), the routine construction of the corresponding Hamiltonian in conjugate momentum variable is not unique. An analogous incertitude is encountered in cosmology models ctv ; nhm , in generalized schemes of Einstein gravity which involve topological invariants, and in theories of higher-curvature gravity ctj .
IV.2 Velocity-independent potentials
@@
$p=p(v)=\left(\frac{1}{4}\right)^{\frac{2}{2k+1}}\frac{1}{(v-1)^{\frac{2}{2k+1}}}\,.$ (33)
Inverting the relation we observe that the velocity variable $v(p)$ emerges as a double-valued function of $p$:
@@
From a classical perspective, in order to avoid an imaginary $v(p)$, one needs to address a nonnegative $p$. This in turn implies that the slope $\frac{\partial L}{\partial v}$ is always positive. It is interesting to note that for the $k=1$ case we are led to the quantum-mechanical supersymmetric structure for the difference $H_{\pm}(x,p)-V(x)$, which reads $p\pm\frac{1}{2\sqrt{p}}$, in the momentum space. The associated spectral properties have been analyzed in the literature ruby ; BGG .
We end our discussion on this example by noting that in the special case where $V(x)=x^{2}$, the branched Hamiltonian is $H_{\pm}(x,p)=x^{2}+U_{\pm}(p)$, wherein it appears as if the roles of the coordinate and the momentum have been interchanged, with $U_{\pm}(p)$ being a momentum-dependent potential that exhibits two branches.
IV.3 Velocity-dependent potentials
@@
Unfortunately, since a Hamiltonian has to be a function of the coordinate and its corresponding canonical momentum, the generality of the form of $H_{\pm}(x,p)$ as derived above is of little use unless we have an explicit inversion of (37) giving $v=v(p)$. We therefore have to go for the specific cases of $k$ and $U(v)$.
IV.3.1 A special case
@@
where $\mu^{3/2}$ has been replaced by $\gamma$. As a final comment, the special case corresponding to $\lambda=0$ and $\gamma=1/4$ conforms to the Hamiltonian (35) advanced in CZ .
V Three more forms of Hamiltonians
V.1 Higher-power Lagrangians
As an extension of (31), the following higher-power Lagrangian was proposed in bspf :
@@ V.1 Higher-power Lagrangians
where we notice that the coefficient $\Lambda$ is non-negative for $0\leq m<\frac{1}{2}$. The main difference from (31) is in the choice of a general function $\sigma(x)$ in place of $\sigma(x)=-1$ as in (31). The other point is that the inverse exponent with respect to the model of Curtright and Zachos CZ has been taken for convenience of calculus. We have omitted the explicit potential function assuming that the interaction re-appears in a more natural manner via a suitable choice of an auxiliary free parameter $\delta$ and that of a nontrivial function $\sigma(x)$. As long as our Lagrangian $L(x,v)$ is of a nonstandard type, we will not feel disturbed by the absence of the explicit potential $V(x)$.
@@ V.1.1 Special case
V.2 Rational-function Lagrangians
In another characteristic example, let us pick up an illustration where $L(x,v)$ is of the reciprocal kind and is defined to be bspf
@@ V.2 Rational-function Lagrangians
It should be remarked that as $\lambda\rightarrow 0$, (53) is just the trial Lagrangian (6), for the choice $\mu(x)=ax^{2}$ and under a suitable identification of the constant parameters. We end by noting that (56) can be re-expressed as a model,
$H_{\pm}(x,p)=\frac{x^{2}}{2m(p)}+U_{\pm}(p),\quad\quad m(p)=\frac{3}{2sp},\quad\quad U_{\pm}(p)=\frac{3}{s}\lambda p\pm 2\sqrt{\frac{p}{s}},$ (57)
with momentum-dependent mass.
V.3 Relativistic free particle
As a final example of dynamics due to non-quadratic Lagrangians, let us re-examine the much-studied problem of a relativistic (free) particle which is described by the Lagrangian
$L(v)=-mc\sqrt{c^{2}-v^{2}},\quad\quad v<c.$ (58)
The conjugate momentum is obtained as
$p=p(v)=\frac{\partial L(v)}{\partial v}=\frac{mcv}{\sqrt{c^{2}-v^{2}}}.$ (59)
This implies that $p^{2}c^{2}-p^{2}v^{2}-m^{2}c^{2}v^{2}=0$. Solving for the velocity gives
$v=v_{\pm}(p)=\pm\frac{pc}{\sqrt{p^{2}+m^{2}c^{2}}}.$ (60)
Thus, the Hamiltonian reads
$H_{\pm}(p)=\frac{c(\pm p^{2}-m^{2}c^{2})}{\sqrt{p^{2}+m^{2}c^{2}}}.$ (61)
In particular, one may consider the ultrarelativistic limit ($v\approx c$) in which $H_{\pm}(p)=\pm pc$. Related to the above example, the reader is referred to Ref. example1 for the example of a spinning particle with the Lagrangian being non-quadratic both in the position and spin variables. Another physically-interesting example is that of an axially-symmetric charged body in an electromagnetic field which is governed by Euler-Poisson equations example2 .
VI Concluding remarks
In the present treatment on the existence of nonstandard Lagrangians, we emphasized, first of all, the existence of certain unusual aspects of their relationship with the associated branched Hamiltonians. Various different examples were discussed; in all of them, the velocity dependence of Lagrangian was not of (homogenous) degree two but contained either powers larger than two or negative powers. This resulted in a nonlinear relationship between the generalized velocity and the conjugate momentum, leading to a multi-valued behavior of the velocity when solved as a function of the momentum (and perhaps the coordinate).
@@ Acknowledgments
Appendix A Jacobi last multiplier
The Jacobi last multiplier whittaker is a very useful tool in the mathematics of dynamical systems. On one hand, given a dynamical system on an $m$-dimensional phase space with $(m-2)$ linearly-independent and known first integrals, it allows one to determine the ‘last’, i.e., $(m-1)$th independent first integral; on the other hand, given a second-order (ordinary) differential equation, it facilitates the computation of a suitable Lagrangian function which describes the dynamics via the Euler-Lagrange equation whittaker ; yan78 ; Leach1 ; Leach2 . Thus, the last multiplier is intimately related with the integrability properties of a dynamical system. Naively, one can define it as follows. Given a vector field $X$ on the phase space, the Jacobi last multiplier $M$ is a factor such that $MX$ has zero divergence. For Hamiltonian systems for which Liouville’s theorem holds, i.e., ${\rm div}\cdot X=0$, the last multiplier is just a constant number as $X$ is already divergence free. On the other hand, $M$ assumes a more non-trivial form if $X$ has non-vanishing divergence, i.e., if ${\rm div}\cdot X\neq 0$, say, for the Liénard equation, which is dissipative and hence, is not volume conserving (see for example, Refs. CG19 ; CG24 ).
Consider a vector field $X$; in local coordinates $\{x_{j}\}$, where $j=1,2,\cdots,m$, one can write its components as $\dot{x}_{j}=X_{j}(\{x_{j}\})$ which define the dynamical system as a system of first order equations. Further, let there be a certain number of first integrals $(F_{1},F_{2},\cdots,F_{k})$, where $k<m$. For any open subset $\Omega\subset\mathbb{R}^{m}$, we define the Jacobi last multiplier to be a function $M:\mathbb{R}^{m}\rightarrow\mathbb{R}$ which is non-negative and defines an invariant measure $\int_{\Omega}Md^{m}x$, i.e.,
[…]
where $\phi_{t}(\Omega)$ is the transformation of the region $\Omega$ under the flow of $X$. Notice that it is necessary that $(F_{1},F_{2},\cdots,F_{k})$ must be independent, ensuring that $dF_{1}\wedge dF_{2}\wedge\cdots\wedge dF_{k}\neq 0$. The above-mentioned invariance condition leads to the equation
$\frac{d}{dt}\Bigg{(}M\frac{\partial(x_{1},x_{2},\cdots,x_{k},x_{k+1},\cdots,x_{m})}{\partial(F_{1},F_{2},\cdots,F_{k},x_{k+1},\cdots,x_{m})}\Bigg{)}=0,$ (63)
which, upon employing the chain rule gives
$\frac{dM}{dt}\frac{\partial(x_{1},x_{2},\cdots,x_{k},x_{k+1},\cdots,x_{m})}{\partial(F_{1},F_{2},\cdots,F_{k},x_{k+1},\cdots,x_{m})}+M\sum_{j=1}^{m}\frac{\partial X_{j}}{\partial x_{j}}\frac{\partial(x_{1},x_{2},\cdots,x_{k},x_{k+1},\cdots,x_{m})}{\partial(F_{1},F_{2},\cdots,F_{k},x_{k+1},\cdots,x_{m})}=0.$ (64)
This is just equivalent to
$\frac{d}{dt}\ln M+\sum_{j=1}^{m}\frac{\partial X_{j}}{\partial x_{j}}=0,$ (65)
which coincides with (11) as appropriate for the system (10). Notice that $\sum_{j=1}^{m}\frac{\partial X_{j}}{\partial x_{j}}$ is just the divergence of $X$ and therefore, $M=1$ or some constant if the vector field has zero divergence.
Now having defined the Jacobi last multiplier, let us demonstrate its use in deriving Lagrangians for a second-order ordinary differential equation with a two-dimensional phase space. Consider the system (10). If it is derivable from the Euler-Lagrange equation
$\frac{d}{dt}\bigg{(}\frac{\partial L(x,\dot{x})}{\partial\dot{x}}\bigg{)}=\frac{\partial L(x,\dot{x})}{\partial x},$ (66)
or,
$\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}\ddot{x}+\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}\partial x}\dot{x}-\frac{\partial L(x,\dot{x})}{\partial x}=0,$ (67)
then, from (10) one should have
$\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}F(x,\dot{x})+\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}\partial x}\dot{x}-\frac{\partial L(x,\dot{x})}{\partial x}=0.$ (68)
Differentiating both sides with respect to $\dot{x}$ gives
$\frac{\partial}{\partial\dot{x}}\bigg{(}\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}F(x,\dot{x})\bigg{)}+\frac{\partial^{3}L(x,\dot{x})}{\partial\dot{x}^{2}\partial x}\dot{x}=0.$ (69)
Defining $\Sigma(x,\dot{x})=\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}$, (69) gives
$\frac{\partial}{\partial\dot{x}}(\Sigma(x,\dot{x})F(x,\dot{x}))+\frac{\partial}{\partial x}(\Sigma(x,\dot{x})\dot{x})=0.$ (70)
But this is just the equation (65) of the last multiplier (if $M$ has no explicit time dependence), i.e., we should identify $\Sigma(x,\dot{x})=M(x,\dot{x})$ and this readily gives (12). See Refs. whittaker ; yan78 ; Leach1 ; Leach2 for more details.